Sam purchased a new winter coat while he was at the LL Bean outlet store. The total cost of the coat including an 6.5% sales tax was $189.57. Find the price of the jacket before tax.
Let x=

Answer by stanbon(75887)

Sam purchased a new winter coat while he was at the LL Bean outlet store. The total cost of the coat including an 6.5% sales tax was $189.57. Find the price of the jacket before tax.
-----
Let "x" be the price before tax:
----
Equation:
x + 0.065x = 189.57
-----
1.065x = 189.57
----
x = 189.57/1.065 =
-----
x = $178.00
=====================
Cheers,
Stan H.